A 76 kg person sits on a 4.0 kg chair. Each leg of the chair makes contact with the floor in a circle that is 1.2 cm in diameter.

Find the pressure exerted on the floor by each leg of the chair, assuming the weight is evenly distributed.
Answer #1

total mass = 76+4 = 80 kg  

mass on each leg = 80/4 = 20 kg

force on each leg = 20*9.8 = 196 N

r = 1.2/2 = 0.6 cm = 0.006 m

Pressure = F/A = 196 / ( pi*0.006^2) =1733020.49 Pa =1.73302049*10^6 Pa on each leg